Patterns within systems of linear equations. Examine the constants in the first equation and describe any patterns. Repeat for the second equation.
x+2y=3
2x-y=-4
How do I determine patterns across the equation and down for the x values?

Answer by stanbon(75887)

Patterns within systems of linear equations. Examine the constants in the first equation and describe any patterns. Repeat for the second equation.
x+2y=3
2x-y=-4
How do I determine patterns across the equation and down for the x values?
---------------------------
"Discribe any patterns" is a very large order.
There are many patterns.
----
To solve the system,
Multiply thru the 1st equation by 2 to get:
2x + 4y = 6
2x - y =-4
------
Subtract the 2nd from the 1st and solve for "y":
5y = 10
y = 2
----
Substitute into 2x-y=-4 and solve for "x":
2x-2 = 4
x-1 = 2
x = 3
---------------
Cheers,
Stan H.